OpenStudy (anonymous):

Graph the first six terms of a sequence where a_1 = -10 and d = 3.

OpenStudy (anonymous):

@ranga

OpenStudy (ranga):

It is an arithmetic sequence with the first term -10 and common difference 3. What are the first six terms? I will write the first 3 and yo do the last three. -10, -10+3 = -7, -7 + 3 = -4, ...

OpenStudy (anonymous):

Can you draw it please.

OpenStudy (ranga):

To get the current term you just add 3 to the previous term. So we have -10, -7, -4, ...

OpenStudy (ranga):

First list the six terms and then we can get to graph it.

OpenStudy (anonymous):

@ranga pls helppp

OpenStudy (anonymous):

@agent0smith

OpenStudy (agent0smith):

"First list the six terms and then we can get to graph it."

OpenStudy (agent0smith):

I'm sure you're capable of adding 3 a few times.

OpenStudy (agent0smith):

"we have -10, -7, -4, ..."